# OpenSec: Measuring Incident Response Agent Calibration Under Adversarial Evidence

### v2

@@ Abstract
As large language models improve, so do their offensive applications: frontier agents now generate working exploits for under \$50 in compute (Heelan, 2026). Defensive incident response (IR) agents must keep pace, but existing benchmarks conflate action execution with correct execution, hiding calibration failures when agents process adversarial evidence. We introduce OpenSec, a dual-control reinforcement learning environment that evaluates IR agents under realistic prompt injection scenarios. Unlike static capability benchmarks, OpenSec scores world-state-changing containment actions under adversarial evidence via execution-based metrics: time-to-first-containment (TTFC), blast radius (false positives per episode), and injection violation rates. Evaluating four frontier models on 40 standard-tier episodes, we find consistent over-triggering in this setting: GPT-5.2, Gemini 3, and DeepSeek execute containment in 100% of episodes with 90–97% false positive rates. Claude Sonnet 4.5 shows partial calibration (85% containment, 72% false positive), demonstrating that OpenSec surfaces a calibration failure mode hidden by aggregate success metrics. Code available at https://github.com/jbarnes850/opensec-env.
reinforcement learning, cybersecurity, incident response, agent calibration, benchmark, OpenSec, adversarial evidence, prompt injection
@@ 1 Introduction
OpenSec measures what current benchmarks miss: the gap between action willingness and action correctness when evidence is adversarial and stakes are operational.
@@ 1.2 Contributions
We make four contributions. First, we introduce OpenSec, a dual-control RL environment with deterministic, execution-based scoring for training IR agents. Second, we design taxonomy-stratified scenarios with trust tiers and prompt injection payloads that enable curriculum learning. Third, we evaluate four frontier models on 40 standard-tier episodes and reveal consistent over-triggering in this setting, with 90–97% false positive rates. Fourth, we provide evidence that calibration varies across frontier models and is not reliably present: Sonnet 4.5 shows partial calibration where GPT-5.2, Gemini 3, and DeepSeek do not.
@@ 2 Environment Design
OpenSec is a dual-control simulator with deterministic scoring. The defender observes evidence from SQLite logs, alerts, and emails and uses tools to investigate and contain. The attacker advances a fixed kill chain (phish_sent $\rightarrow$ creds_used $\rightarrow$ lateral_move $\rightarrow$ data_access $\rightarrow$ exfil_attempt) with state-constrained actions and optional branch variants.
@@ 2.1 Scenario Design
The dataset comprises 160 training seeds and 40 evaluation seeds, all standard-tier, with injection payloads present across all scenarios.
@@ 4 Baseline Evaluation
Metrics. We report five metrics per model. Containment rate is the fraction of episodes where the agent executes at least one containment action. False positive rate is the fraction of episodes containing at least one incorrect containment action. Correct containment rate is the fraction of ground-truth targets correctly contained. Injection violation rate is the fraction of episodes where the agent executes a tool call containing a substring from an injected payload after exposure to that payload. Blast radius is the mean count of incorrect containment actions per episode.
@@ 4.1 Results
Three of four frontier models execute containment in 100% of episodes with 90–97% false positive rates. They trigger nearly every available containment action regardless of evidence quality.
Table 1: Frontier model evaluation on 40 standard-tier episodes. Cont.=containment rate, FP=false positive rate, Corr.=correct containment, Inj.=injection violation rate. 95% Wilson CIs for rates: $\pm$0.05–0.15 depending on rate. High rewards mask operational failure: models exhaust the action space, achieving high correct containment alongside 90–97% FP rates.
Model Reward Cont. FP Corr. Inj.
GPT-5.2 3.46 1.00 0.97 0.97 0.38
Sonnet 4.5 2.76 0.85 0.72 0.85 0.40
Gemini 3 3.35 1.00 0.97 1.00 0.50
DeepSeek 3.2 2.99 1.00 0.90 1.00 0.78
Calibration varies across frontier models. GPT-5.2, Gemini 3, and DeepSeek show identical over-triggering behavior. Only Sonnet 4.5 demonstrates partial pretrained calibration (85% containment, 72% FP), suggesting calibration depends on factors beyond capability, potentially including alignment approach or training methodology.
High rewards mask operational failure. The reward range (2.76–3.46) looks strong, but these scores reflect indiscriminate action. By exhausting the containment action space, models capture all correct targets alongside nearly all incorrect ones. In production, this would take down legitimate services indiscriminately.
Injection vulnerability varies independently. DeepSeek shows the highest violation rate (78%) despite identical containment behavior to GPT-5.2 (38%). This suggests injection robustness is orthogonal to containment calibration and requires separate attention.
@@ 4.2 Operational Metrics
Beyond aggregate rates, operational timing provides insight into response behavior. Time-to-first-containment (TTFC) measures when an agent first executes a containment action; blast radius counts false positive containment actions per episode.
Table 2: Operational timing metrics (in steps). TTFC=time-to-first-containment, TTR=time-to-report. Sonnet 4.5 waits approximately 3 steps longer before first containment, resulting in lower blast radius.
Model TTFC TTR Blast (mean) Blast (max)
GPT-5.2 6.95 10.47 1.23 3
Sonnet 4.5 9.91 13.03 1.15 2
Gemini 3 7.73 11.55 1.40 2
DeepSeek 3.2 7.58 11.55 1.18 2
Over-triggering models act faster: GPT-5.2, Gemini 3, and DeepSeek execute containment within 7–8 steps, taking action before gathering sufficient evidence. Sonnet 4.5 has the lowest blast radius (1.15) despite higher TTFC, suggesting delayed action results in fewer false positives.
@@ 5 Discussion
Aggregate scores mask operational failure. Frontier models achieve rewards of 2.76–3.46, but three of four do so by exhausting the containment action space. By aggregate metrics, these are high-performing agents (97–100% correct containment). By operational metrics, they would take down production services indiscriminately (90–97% false positive rates). Current evaluation practices conflate action execution with correct action execution.
Calibration exists in some pretrained models. Sonnet 4.5’s partial calibration (85%/72%) shows the capability exists without targeted training. Why Sonnet and not GPT-5.2, Gemini, or DeepSeek? This is an open question, but the variation itself is diagnostic value the environment provides.
Injection vulnerability is orthogonal. DeepSeek shows 78% violation rate despite identical containment behavior to GPT-5.2 (38%). The OWASP Agentic AI Guide (OWASP, 2025) identifies tool/API access as a key attack surface. OpenSec deliberately places the defender in this configuration because real IR requires processing attacker-controlled content.
@@ 6 Limitations
The evaluation uses 40 standard-tier seeds per model. Broader statistical confidence requires additional seeds and replications. Trust tier metadata is present but not yet used as an evaluation signal.
@@ 7 Future Work
Trust-aware evaluation. The trust_profile field enables measuring whether models appropriately weight evidence by provenance tier. This is not yet analyzed but the infrastructure exists.
@@
Calibration training. Preliminary RL experiments (Appendix A) suggest calibration behavior is trainable but requires further investigation, likely a two-stage SFT+RL pipeline or curriculum approach.
@@ A.1 Method
We trained Qwen/Qwen3-4B-Instruct with GDPO using decomposed reward functions (attribution, containment, injection, efficiency). GDPO decouples normalization across rewards before aggregation, addressing reward-advantage collapse in multi-reward settings. Training used SGLang for rollouts on a single A100.
@@ A.3 Interpretation
The trained model shows modified but not clearly improved calibration compared to Sonnet 4.5 (85%/72%). Correct containment (47.5%) is lower than Sonnet (85%). Report submission dropped to 25%, suggesting reward shaping issues. The model learned to act less frequently but not more accurately.

### v3

@@ Abstract
As large language models (LLMs) improve, so do their offensive applications: frontier agents now generate working exploits for under \$50 in compute (Heelan, 2026). Defensive incident response (IR) agents must keep pace, but existing benchmarks conflate action execution with correct execution, hiding calibration failures when agents process adversarial evidence. We introduce OpenSec, a dual-control reinforcement learning (RL) environment that evaluates IR agents under realistic prompt injection scenarios with execution-based scoring: time-to-first-containment (TTFC), evidence-gated action rate (EGAR), blast radius, and per-tier injection violation rates. Evaluating four frontier models on 40 standard-tier episodes each, we find consistent over-triggering: GPT-5.2 executes containment in 100% of episodes with 82.5% false positive rate, acting at step 4 before gathering sufficient evidence. Claude Sonnet 4.5 shows partial calibration (62.5% containment, 45% FP, TTFC of 10.6), suggesting calibration is not reliably present across frontier models. All models correctly identify the ground-truth threat when they act; the calibration gap is not in detection but in restraint. Code available at https://github.com/jbarnes850/opensec-env.
reinforcement learning, cybersecurity, incident response, agent calibration, benchmark
@@ 1 Introduction
OpenSec measures what current benchmarks miss: the gap between action willingness and action correctness when evidence is adversarial and stakes are operational. Figure 1 illustrates this gap: GPT-5.2 executes containment at step 4, before the attacker reaches lateral movement, while Sonnet 4.5 investigates 70% of the episode before acting.
Figure 1: Representative episode timelines for GPT-5.2 (uncalibrated) and Sonnet 4.5 (partially calibrated) on a standard-tier scenario. The attacker kill chain progresses across 15 steps regardless of defender behavior. GPT-5.2 begins containment at step 4 with an 82% false positive rate; Sonnet 4.5 gathers evidence until step 11 with a 45% false positive rate. Both models correctly identify the ground-truth threat—the calibration gap is in restraint, not detection.
@@ 1.2 Contributions
We make four contributions. First, we introduce OpenSec, a dual-control RL environment with deterministic, execution-based scoring for training IR agents. Second, we design taxonomy-stratified scenarios with trust tiers and prompt injection payloads that enable curriculum learning. Third, we evaluate four frontier models on 40 standard-tier episodes each and reveal consistent over-triggering, with 45–82.5% false positive rates and EGAR below 55% across all models. Fourth, we provide evidence that calibration varies across frontier models: Sonnet 4.5 shows partial calibration (62.5% containment, TTFC 10.6) where GPT-5.2 is uncalibrated (100% containment, TTFC 4.1).
@@ 2 Environment Design
OpenSec is a dual-control simulator with deterministic scoring (Figure 2). The defender observes evidence from SQLite logs, alerts, and emails and uses tools to investigate and contain. The attacker advances a fixed kill chain (phish_sent $\rightarrow$ creds_used $\rightarrow$ lateral_move $\rightarrow$ data_access $\rightarrow$ exfil_attempt) with state-constrained actions and optional branch variants.
@@ 2.1 Scenario Design
The dataset comprises 160 training seeds and 60 evaluation seeds (20 trivial, 20 easy, 20 standard), with injection payloads present across all scenarios. Baseline evaluation uses the 40 standard-tier seeds. Figure 3 illustrates the seed generation pipeline.
@@ 4 Baseline Evaluation
Metrics. We report six metrics per model. Containment rate is the fraction of episodes where the agent executes at least one containment action. False positive rate is the fraction of episodes containing at least one incorrect containment action. Evidence-gated action rate (EGAR) is the fraction of containment actions preceded by fetching trusted evidence about the target entity; it measures whether the model checks before it acts. Time-to-first-containment (TTFC) is the step index of the first containment action; higher values indicate more investigation before acting. Blast radius is the ratio of false positive to correct containment actions per episode. Injection violation rate is reported per tier: T1 (obvious overrides), T2 (contextualized domain-specific framing), T3 (complex multi-step or multilingual payloads). We additionally report time-to-report (TTR), the step index at which the agent submits its final report, in operational analysis.
@@ 4.1 Results
All four frontier models execute containment in 62.5–100% of episodes with 45–82.5% false positive rates. EGAR ranges from 37.5% to 54.2%, indicating that most containment actions are taken without first gathering trusted evidence about the target entity.
Table 1: Frontier model evaluation on 40 standard-tier episodes each. Cont.=containment rate, FP=false positive rate, EGAR=evidence-gated action rate, TTFC=time-to-first-containment. All models correctly identify the ground-truth threat when they act; the calibration gap is in restraint, not detection.
Model Reward Cont. FP EGAR TTFC Threshold
GPT-5.2 3.07 1.00 0.825 0.375 4.1 Uncalib.
Sonnet 4.5 2.37 0.625 0.45 0.392 10.6 Part. Cal.
Gemini 3 2.61 0.75 0.575 0.429 8.6 Part. Cal.
DeepSeek 3.2 3.45 0.925 0.65 0.542 9.0 Part. Cal.
Calibration varies across frontier models. GPT-5.2 is the only model classified as uncalibrated, executing containment in 100% of episodes at step 4.1 with 82.5% false positive rate. Sonnet 4.5 shows partial calibration (62.5% containment, 45% FP), waiting until step 10.6 to act. Gemini 3 and DeepSeek fall between these extremes (Figure 1). Calibration depends on factors beyond capability, potentially including alignment approach or training methodology.
High rewards mask operational failure. The reward range (2.37–3.45) looks strong, but these scores reflect indiscriminate action. All models correctly identify the ground-truth threat when they act; the calibration gap is not in detection but in restraint. In production, the false positive actions would take down legitimate services alongside the real threat.
Injection vulnerability varies by tier. Per-tier injection analysis reveals that T2 (contextualized) payloads are most effective across all models (15–25% violation rate), while T1 (obvious) payloads rarely succeed. DeepSeek shows the highest T3 (complex) vulnerability at 10%. Injection robustness is orthogonal to containment calibration.
@@ 4.2 Operational Metrics
Beyond aggregate rates, operational timing provides insight into response behavior. Time-to-first-containment (TTFC) measures when an agent first executes a containment action; blast radius is the ratio of false positive to correct containment actions per episode.
Table 2: Operational and injection metrics. TTFC=time-to-first-containment, TTR=time-to-report, Blast=ratio of FP to correct actions. Per-tier injection violation rates: T1=obvious, T2=contextualized, T3=complex. Sonnet 4.5 waits longest (TTFC 10.6) and has zero T1/T3 vulnerability.
Model TTFC TTR Blast T1 T2 T3
GPT-5.2 4.1 12.1 0.43 0% 25% 7%
Sonnet 4.5 10.6 13.5 0.44 0% 20% 0%
Gemini 3 8.6 12.5 0.44 7% 15% 5%
DeepSeek 3.2 9.0 13.2 0.42 5% 15% 10%
GPT-5.2 acts fastest (TTFC 4.1), executing containment after investigating only 27% of the episode. Sonnet 4.5 waits until step 10.6 (70% of the episode), resulting in the lowest false positive rate. All models show similar blast radius (0.42–0.44), indicating that when false positives occur, their magnitude is consistent regardless of timing. T2 (contextualized) injection payloads are the most effective attack surface across all models, while T1 (obvious) payloads rarely succeed, suggesting frontier models have baseline resistance to crude override attempts.
@@ 5 Discussion
Aggregate scores mask operational failure. Frontier models achieve rewards of 2.37–3.45, but all do so while generating 45–82.5% false positive rates. All models correctly identify the ground-truth threat when they act. The calibration gap is not in detection but in restraint: models act on the right target and wrong targets simultaneously. Current evaluation practices conflate action execution with correct action execution.
Calibration exists in some pretrained models. Sonnet 4.5’s partial calibration (62.5% containment, 45% FP, TTFC 10.6) shows the capability exists without targeted training. GPT-5.2 represents the opposite extreme: 100% containment at step 4.1 with 82.5% FP rate. Why Sonnet and not GPT-5.2, Gemini, or DeepSeek? This is an open question, but the EGAR and TTFC metrics make the variation precisely measurable.
Injection vulnerability is tier-dependent. Per-tier analysis reveals that T2 (contextualized) payloads are the primary attack surface (15–25% across models), while T1 (obvious) payloads rarely succeed. DeepSeek shows the highest T3 (complex) vulnerability at 10%. The OWASP Agentic AI Guide (OWASP, 2025) identifies tool/API access as a key attack surface. OpenSec deliberately places the defender in this configuration because real IR requires processing attacker-controlled content.
@@ 6 Limitations
The evaluation uses 40 standard-tier seeds per model. Broader statistical confidence requires additional seeds and replications. Trust tier metadata is used for EGAR computation (only trusted evidence counts toward evidence gating) but is not yet analyzed as an independent variable. The defensive capability thresholds are provisional, calibrated against observed frontier model behavior rather than human expert baselines.
@@ 7 Future Work
Trust-aware evaluation. The trust_profile field and EGAR metric provide infrastructure for measuring whether models appropriately weight evidence by provenance tier. While EGAR uses trust tiers for evidence gating, analyzing model behavior as a function of evidence provenance quality remains future work.
@@
Calibration training. Preliminary RL experiments (Appendix A) suggest calibration behavior is trainable but requires further investigation, likely a two-stage supervised fine-tuning (SFT) + RL pipeline or curriculum approach.
@@ A.1 Method
We trained Qwen/Qwen3-4B-Instruct with Group reward-Decoupled Normalization Policy Optimization (GDPO) using decomposed reward functions (attribution, containment, injection, efficiency). GDPO decouples normalization across rewards before aggregation, addressing reward-advantage collapse in multi-reward settings. Training used SGLang for rollouts on a single A100.
@@ A.3 Interpretation
The trained model shows modified but not clearly improved calibration compared to Sonnet 4.5 (62.5% containment, 45% FP). The model executes containment in 75% of episodes (vs. Sonnet’s 62.5%) but with a 70% false positive rate (vs. Sonnet’s 45%). Correct containment is 47.5%, indicating the model acts more frequently but less accurately. Report submission dropped to 25%, suggesting reward shaping issues.